In a market with Pd = 150 – 3 Qd and Ps = 50 + Qs , the government has regulated maximum of 20 units and is concerned if this move was maybe a little too harsh, especially if the deadweight loss to consumers (voters) is greater than the deadweight loss to producers
1
Expert's answer
2020-12-23T04:32:50-0500

If Pd = 150 – 3Qd and Ps = 50 + Qs, then in equilibrium:

"150 \u2013 3Q = 50 + Q,"

4Q = 100,

Q = 25 units.

P = 50 + 25 = 75.

the government has regulated maximum of 20 units, then Pd = 150 - 3×20 = 90 and Ps = 50 + 20 = 70.

The deadweight loss to consumers (voters) is:

"DWLc = 0.5\u00d7(25 - 20)\u00d7(90 - 75) = 37.5,"

And the deadweight loss to producers is:

"DWLp = 0.5\u00d7(25 - 20)\u00d7(75 - 70) = 12.5."